A,B, and C form a sequence of two digit prime numbers less than 100.The difference between A and B; B and C are equal.Find the value of B.
Answer by Edwin McCravy(20054)

There are 27 different situations where that is true,
but some use the same values of B, which is what you
are looking for:

1. A = 11 B = 17 C = 23  B - A = 6 and C - B also = 6
2. A = 11 B = 29 C = 47  B - A = 18 and C - B also = 18
3. A = 11 B = 41 C = 71  B - A = 30 and C - B also = 30
4. A = 11 B = 47 C = 83  B - A = 36 and C - B also = 36
5. A = 13 B = 37 C = 61  B - A = 24 and C - B also = 24
6. A = 13 B = 43 C = 73  B - A = 30 and C - B also = 30
7. A = 17 B = 23 C = 29  B - A = 6 and C - B also = 6
8. A = 17 B = 29 C = 41  B - A = 12 and C - B also = 12
9. A = 17 B = 53 C = 89  B - A = 36 and C - B also = 36
10. A = 19 B = 31 C = 43  B - A = 12 and C - B also = 12
11. A = 19 B = 43 C = 67  B - A = 24 and C - B also = 24
12. A = 23 B = 41 C = 59  B - A = 18 and C - B also = 18
13. A = 23 B = 47 C = 71  B - A = 24 and C - B also = 24
14. A = 23 B = 53 C = 83  B - A = 30 and C - B also = 30
15. A = 29 B = 41 C = 53  B - A = 12 and C - B also = 12
16. A = 29 B = 59 C = 89  B - A = 30 and C - B also = 30
17. A = 31 B = 37 C = 43  B - A = 6 and C - B also = 6
18. A = 37 B = 67 C = 97  B - A = 30 and C - B also = 30
19. A = 41 B = 47 C = 53  B - A = 6 and C - B also = 6
20. A = 43 B = 61 C = 79  B - A = 18 and C - B also = 18
21. A = 47 B = 53 C = 59  B - A = 6 and C - B also = 6
22. A = 47 B = 59 C = 71  B - A = 12 and C - B also = 12
23. A = 53 B = 71 C = 89  B - A = 18 and C - B also = 18
24. A = 59 B = 71 C = 83  B - A = 12 and C - B also = 12
25. A = 61 B = 67 C = 73  B - A = 6 and C - B also = 6
26. A = 61 B = 79 C = 97  B - A = 18 and C - B also = 18
27. A = 67 B = 73 C = 79  B - A = 6 and C - B also = 6

Edwin